OpenStudy (anonymous):

The ninth grade class at a local high school needs to purchase a park permit for 250.00 for their upcoming class picnic. Eah ninth grader attending the picnic pays 0.75. Each guest pays 1.25. If 200 ninth graders attend the picnic, write an inequality that can be used to determine the number of guests, x, needed to cover the cost of the permit?

OpenStudy (anonymous):

ok so . this is simple

OpenStudy (anonymous):

ninth grader=x guest =y

OpenStudy (anonymous):

0.75x+1.25y>=250

OpenStudy (anonymous):

since there are 200 ninth graders you know that 0.75(200)+1.25y>=250

OpenStudy (anonymous):

150+1.25y>=250 this is your answer. or 1.25y>=100 depends on if you have choices or not